In the experiment determined the value of magnetic induction, the

electrons are accelerated from a stationary state with a potential difference

350V. Electrons move in circles. Know the radius

The measured trajectory is 7.5 cm and the magnetic induction is perpendicular to the beam

electron. (a) What is the value of induction? (b) Determination

the angular velocity of the electron.


1
Expert's answer
2021-01-24T14:24:00-0500

"2\pi r=7.5\to r=1.19(cm)"


(a) "qvB=m\frac{v^2}{r}\to B=\frac{mv}{qr}"


"qV=mv^2/2\to v=\sqrt{2qV/m}="


"=\sqrt{2\cdot 1.6\cdot 10^{-19}\cdot 350/9.1\cdot10^{-31}}=11.1\cdot10^6(m/s)" .


"B=\frac{9.1\cdot 10^{-31}\cdot11.1\cdot10^6}{1.6\cdot10^{-19}\cdot 1.19\cdot10^{-2}}=5.3(mT)" . Answer


(b) "\omega=v/r=11.1\cdot10^6/0.0119=933\cdot10^6(rad/s)" . Answer